# Niche Dynamics in Complex Online Community Ecosystems

### v1

@@ Abstract
Online communities are important organizational forms where members socialize and share information. Curiously, different online communities often overlap considerably in topic and membership. Recent research has investigated competition and mutualism among overlapping online communities through the lens of organizational ecology; however, it has not accounted for how the nonlinear dynamics of online attention may lead to episodic competition and mutualism. Neither has it explored the origins of competition and mutualism in the processes by which online communities select or adapt to their niches. This paper presents a large-scale study of 8,806 Reddit communities belonging to 1,919 clusters of high user overlap over a 5-year period. The method uses nonlinear time series methods to infer bursty, often short-lived ecological dynamics. Results reveal that mutualism episodes are longer lived and slightly more frequent than competition episodes. Next, it tests whether online communities find their niches by specializing to avoid competition using panel regression models. It finds that competitive ecological interactions lead to decreasing topic and user overlaps; however, changes that decrease such niche overlaps do not lead to mutualism. The discussion proposes that future designs may enable online community ecosystem management by informing online community leaders to organize “spin-off” communities or via feeds and recommendations.
@@ Introduction
Interdisciplinary scholarship has explored these questions through the lens of organizational ecology (TeBlunthuis and Hill 2022; Zhu, Kraut, and Kittur 2014; Zhu et al. 2014; Wang, Butler, and Ren 2012). Organizational ecology considers the role of environmental and relational forces in the development of organizations and industries (Hannan and Freeman 1989; McPherson 1983; Hawley 1986; Aldrich and Ruef 2006; van de Ven and Poole 1995). Early studies in this vein used density dependence theory to relate the overlaps in community membership to growth and survival (Zhu, Kraut, and Kittur 2014; Zhu et al. 2014; Wang, Butler, and Ren 2012). Recent work by TeBlunthuis and Hill (2022) introduces an alternative approach using time-series analysis to directly infer competitive and mutualistic relationships; however, their analysis is based on an unrealistic assumption in the online community context. They modeled intercommunity ecological relationships as static over time. However, the dynamics of attention online are often bursty (Ratkiewicz et al. 2010).
@@
This work contributes to the social science of online communities by (1) inferring time-varying nonlinear ecological dynamics, (2) shedding new light on findings from prior work, (3) engaging the organizational ecology concepts of structural intertia and competitive exclusion to theorize the role of adaptation in producing the widespread mutualism observed in prior studies and, (4) testing the resulting hypotheses to (5) synthesize insights for leaders and designers of online communities. The nonlinear time-series analysis reveals that ecological relationships tend to happen in bursty episodes and that mutualism has similar strength to competition; however mutualism happens more often and for longer durations. The panel models of niche adaptation found that online communities in competition tend to reduce their overlaps; however such changes lead to increases, not decreases in competition. The latter result is consistent with previous research findings that online communities tend to become oligarchical and rigid (TeBlunthuis, Shaw, and Hill 2018; Hill and Shaw 2019), and that organizational change often has negative outcomes (Hannan and Freeman 1989). Online community designers and leaders can attempt to manage online ecosystems by filling ecological roles. Feed or recommendation algorithms can be designed to adapt existing communities’ niches by purposfully nudging individuals. That said, given the challenges of organizational change, a more pragmatic approach is likely to building new communities as “spin-offs” or “seeds” from existing ones.
@@
Recent scholarship has demonstrated that interdependence among online communities is widespread, important to explaining success and failure, and likely to provide new insights for designing and managing communities (Cunha et al. 2019; Mitts, Pisharody, and Shapiro 2022; Chandrasekharan et al. 2017; Kairam, Wang, and Leskovec 2012; Tan and Lee 2015; Tan 2018; TeBlunthuis et al. 2022; Vincent, Johnson, and Hecht 2018). Most relevant to this study are studies that adopt the theoretical lens of organizational ecology to understand competition and mutualism between overlapping online communities (TeBlunthuis and Hill 2022; Wang, Butler, and Ren 2012; Zhu, Kraut, and Kittur 2014; Zhu et al. 2014). The initial work in this line emphasizes competition and finds that overlapping Usenet groups are likely to compete (Wang, Butler, and Ren 2012). Similarly, in chapter 6 of the prominent textbook Building Successful Online Communities, Resnick et al. (2012) recommend that new online communities “carve out a useful and defendable niche in the ecology of competing communities.”
@@
Even without capacities for top-down structural change, members may tend to contribute to communities in ways that lead to specialization (Van Koevering, Ye, and Kleinberg 2024). Online communities are typically “open” organizations where individuals can freely participate in multiple communities at once and share or repost the same content in different communities (Butler and Wang 2011). When some members become unhappy, an online community characterized by structural inertia is unlikely to change policies or leadership to satisfy them. However, in a system of overlapping open communities, unhappy members can “exit” and migrate to a second overlapping community and bring their contributions along (Frey and Sumner 2018; Hirschman 1970).
@@
The competitive exclusion principle predicts that specialization reduces competition over rival resources between overlapping online communities (Armstrong and McGehee 1980; Hardin 1960). Resource partitioning theory and niche width theory in organizational ecology explain the relationship between competition and overlap (Baum and Shipilov 2006; Dobrev, Kim, and Carroll 2003; Dobrev, Kim, and Hannan 2001; Carroll 1985). Sometimes, organizations can adapt to avoid competition as when automotive manufacturers with greater niche overlaps undergo greater transformations during periods of technological change (Dobrev, Kim, and Carroll 2003). Such change can be risky; however increasing specialization to avoid competition can be necessary for organizational survival (Baum and Shipilov 2006). Importantly, the ecological interactions between two online communities can involve a mixture of mutualistic and competitive components. For instance, two communities may compete over contributors’ time and effort while simultaneously benefiting from sharing information or content. Even if the relationship between two communities is mutualistic overall, a decrease in niche overlap can reduce the competitive component of the relationship and increase the overall mutualism. Therefore, the final two hypotheses test whether increases (decreases) in niche overlaps will lead to subsequent (decreases) increases in (mutualism) competition.
@@
The analysis proceeds as follows and summarized in Figure 1: First, a clustering pipeline discovers groups of highly overlapping subreddits based on membership similarity. Then, the S-Map algorithm measures competition and mutualism episodes with which to test H1. Finally, the panel regression models test H2 and H3 using weekly similarity measures based on membership and content and competition and mutualism that the S-Map measured. The next section documents the measures and analytic plan, and the results follow.
@@
[…]
@@
The units of analysis in all of the regression models are dyadic relationships between subreddits. This dyadic structure violates the assumption that observations are drawn independently from an identical distribution. For instance, when a subreddit’s topic changes, resulting changes in its overlaps in all related subreddits will be correlated. Therefore, the analysis uses the “cluster–robust variance estimation for dyadic data” technique developed by Aronow, Samii, and Assenova and implemented in the dyadRobust R package [^1] https://rdrr.io/github/jbisbee1/dyadRobust/ that provides consistent estimates of standard errors for dyadic data with repeated measures (Aronow, Samii, and Assenova 2015). The Hypotheses are tested via 95% confidence intervals calculated by multiplying the standard errors by 1.96. The dataset is sufficiently large to allow normal distribution approximation.
@@ Discussion
Similar to ecologists caretaking natural environments, designers and leaders of online communities may seek to intervene to manage their ecosystems. They could introduce competitors of problematic communities or develop mutualistic communities to complement existing ones. Should they do so by building a new community or by adapting an existing one? Building new communities is notoriously difficult, prone to failure, and seems slow compared to the timescales of mutualism and competition (Schweik and English 2012); therefore, adaptation might seem more practicable. However, this study’s findings suggest that communities face structural intertia and may struggle to adapt mutualistically; still, solutions are conceivable. When existing communities spawn new communities that have strong connections with their parents, the new communities can grow faster (Tan 2018). Organizing a community’ members to “spin-off” or “seed” a new community may thus be a viable strategy to fill needed roles in online commuity ecosystems. The “parent” community should create a channel where members can collaborate to design the new community and coordinate their efforts to build it. Of course, doing so is only possible through the efforts of interested members.
@@
A second set of threats emerges from the panel regressions. Although fixed effect panel models are consistent even in the presence of time-constant omitted variables (Wooldridge 2011), time-varying unobserved variables, such as trends in topics’ popularity, can potentially bias estimates. Therefore, readers should not draw causal interpretations from these models. In general, this study reveals the overall trends of competition, mutualism, and overlaps on Reddit at a large scale. However, the wide view and analytical approach obscure the particular mechanisms driving competition and mutualism among these communities and when they are likely to arise. Future fine-grained case studies and experiments should seek to uncover such mechanisms.
@@ Acknowledgments
Text from a draft of this article was included as part of the author’s PhD dissertation at the Department of Communication at the University of Washington. He thanks his committee—Professors Benjamin Mako Hill, Kirsten Foot, Aaron Shaw, David McDonald and Emma Spiro, for their generous support, wise advice and insightful comments. Versions of this paper received helpful feedback at the International Communication Association’s 2022 annual meeting and at the International Conference for Computational Social Science. Special thanks to Simone Cenci for advice on the regularized S-MAP. Thanks to the Community Data Science Collective for additional feedback and support.
@@
Ethics statement
This work’s mode of inquiry is to analyze large-scale data from social media that is publicly available. Some have concerns about the ethics of such analysis, such as the possibility that the people whose activities created the data we analyze would have been likely to expect their activity would be so used. The author is sympathetic to these serious concerns. That said, the nature of this analysis exposes no individual to scrutiny and our reporting obscures even the identities of the communities we analyze. As a result, the likely and potential harms of this work to individuals are low.

### v2

@@ Abstract
Online communities are important organizational forms where members socialize and share information. Curiously, different online communities often overlap considerably in topic and membership. Recent research has investigated competition and mutualism among overlapping online communities through the lens of organizational ecology; however, it has not accounted for how the nonlinear dynamics of online attention may lead to episodic competition and mutualism. Neither has it explored the origins of competition and mutualism in the processes by which online communities select or adapt to their niches. This paper presents a large-scale study of 8,806 Reddit communities belonging to 1,919 clusters of high user overlap over a 5-year period. The method uses nonlinear time series methods to infer bursty, often short-lived ecological dynamics. Results reveal that mutualism episodes are longer lived and slightly more frequent than competition episodes. Next, it tests whether online communities find their niches by specializing to avoid competition using panel regression models. It finds that competitive ecological interactions lead to decreasing topic and user overlaps; however, changes that decrease such niche overlaps do not lead to mutualism. The discussion considers future designs for online community ecosystem management.
@@ Introduction
Interdisciplinary scholarship has explored these questions through the lens of organizational ecology (TeBlunthuis and Hill 2022; Zhu, Kraut, and Kittur 2014; Zhu et al. 2014; Wang, Butler, and Ren 2012). Organizational ecology considers the role of environmental and relational forces in the development of organizations and industries (Hannan and Freeman 1989; McPherson 1983; Hawley 1986; Aldrich and Ruef 2006; van de Ven and Poole 1995). Early studies in this vein used density dependence theory to relate the overlaps in community membership to growth and survival (Zhu, Kraut, and Kittur 2014; Zhu et al. 2014; Wang, Butler, and Ren 2012). Recently, TeBlunthuis and Hill (2022) introduced an alternative approach using time-series analysis to directly infer competitive and mutualistic relationships; however, their analysis is based on an unrealistic assumption in the online community context. They modeled intercommunity ecological relationships as static over time. However, the dynamics of attention online are often bursty (Ratkiewicz et al. 2010).
@@
This work contributes to the social science of online communities by (1) inferring time-varying nonlinear ecological dynamics, (2) shedding new light on findings from prior work, (3) engaging the organizational ecology concepts of structural intertia and competitive exclusion to theorize the role of adaptation in producing the widespread mutualism observed in prior studies and, (4) testing the resulting hypotheses to (5) synthesize insights for leaders and designers of online communities. The nonlinear time-series analysis reveals that ecological relationships tend to happen in bursty episodes and that mutualism has similar strength to competition; however mutualism happens more often and for longer durations. The panel models of niche adaptation found that online communities in competition tend to reduce their overlaps; however such changes lead to increases, not decreases in competition. The latter result is consistent with previous research findings that online communities tend to become oligarchical and rigid (TeBlunthuis, Shaw, and Hill 2018; Hill and Shaw 2019), and that organizational change often has negative outcomes (Hannan and Freeman 1989). Online community designers and leaders can attempt to manage online ecosystems by filling ecological roles. Feed or recommendation algorithms can be designed to adapt existing communities’ niches by purposfully nudging individuals. That said, given the challenges of organizational change, a more pragmatic approach is likely to building new communities as spin-offs or offshoots from existing ones.
@@
Recent scholarship has demonstrated that interdependence among online communities is widespread, important to explaining success and failure, and likely to provide new insights for designing and managing communities (Cunha et al. 2019; Mitts, Pisharody, and Shapiro 2022; Chandrasekharan et al. 2017; Kairam, Wang, and Leskovec 2012; Tan and Lee 2015; Tan 2018; TeBlunthuis et al. 2022; Vincent, Johnson, and Hecht 2018). Most pertinent are studies that adopt the theoretical lens of organizational ecology to understand competition and mutualism between overlapping online communities (TeBlunthuis and Hill 2022; Wang, Butler, and Ren 2012; Zhu, Kraut, and Kittur 2014; Zhu et al. 2014). The initial work in this line emphasizes competition and finds that overlapping Usenet groups are likely to compete (Wang, Butler, and Ren 2012). Similarly, in chapter 6 of the prominent textbook Building Successful Online Communities, Resnick et al. (2012) recommend that new online communities “carve out a useful and defendable niche in the ecology of competing communities.”
@@
Even without capacities for top-down structural change, members may tend to contribute to communities in ways that lead to specialization (Van Koevering, Ye, and Kleinberg 2024). Online communities are typically “open” organizations where individuals can freely participate in multiple communities at once and share or repost the same content in different communities (Butler and Wang 2011). When some members become unhappy, an online community characterized by structural inertia is unlikely to change policies or leadership to satisfy them. However, in a system of overlapping open communities, unhappy members can “exit” and migrate to an overlapping community and bring their contributions along (Frey and Sumner 2018; Hirschman 1970).
@@
The competitive exclusion principle predicts that specialization reduces competition over rival resources (Armstrong and McGehee 1980; Hardin 1960). Resource partitioning theory and niche width theory in organizational ecology explain the relationship between competition and overlap (Dobrev, Kim, and Carroll 2003; Dobrev, Kim, and Hannan 2001; Carroll 1985). Sometimes, organizations can adapt to avoid competition as when automotive manufacturers with greater niche overlaps undergo greater transformations during periods of technological change (Dobrev, Kim, and Carroll 2003). Such change can be risky; however increasing specialization to avoid competition can be necessary for organizational survival (Baum and Shipilov 2006). Importantly, the ecological interactions between two online communities can involve a mixture of mutualistic and competitive components. For instance, two communities may compete over contributors’ time and effort while simultaneously benefiting from sharing information or content. Even if the relationship between two communities is mutualistic overall, a decrease in niche overlap can reduce the competitive component of the relationship and increase the overall mutualism. Therefore, the final two hypotheses test whether increases (decreases) in niche overlaps will lead to subsequent (decreases) increases in (mutualism) competition.
@@
The analysis proceeds as follows and summarized in Figure 1: First, a clustering pipeline discovers groups of highly overlapping subreddits based on membership similarity. The S-Map algorithm then measures competition and mutualism episodes with which to test H1. Finally, the panel regression models test H2 and H3 using weekly similarity measures based on membership and content and competition and mutualism that the S-Map measured. The next section documents the measures and analytic plan, and the results follow.
@@
Weekly topic overlap $T_{i,j,t}$ between subreddits is similar to weekly user overlap, but using token-frequency inverse-document-frequency (TF-IDF) vectors instead of author-frequency vectors. Pointwise mutual information (PWMI) detects phrases for inclusion in the TF-IDF vectors. Phrases of up to four terms were used as it was the maximum computationally tractable given the available resources. The number of possible phrases grows exponentially with the phrase length.
@@
The units of analysis in all of the regression models are dyadic relationships between subreddits. This dyadic structure violates the assumption that observations are drawn independently from an identical distribution. When a subreddit’s topic changes, for instance, resulting changes in its overlaps in all related subreddits will be correlated. Therefore, the analysis uses the “cluster–robust variance estimation for dyadic data” technique developed by Aronow, Samii, and Assenova and implemented in the dyadRobust R package [^1] https://rdrr.io/github/jbisbee1/dyadRobust/ that provides consistent estimates of standard errors for dyadic data with repeated measures (Aronow, Samii, and Assenova 2015). The Hypotheses are tested via 95% confidence intervals calculated by multiplying the standard errors by 1.96. The dataset is sufficiently large to allow normal distribution approximation.
@@ Discussion
Similar to ecologists caretaking natural environments, designers and leaders of online communities may seek to intervene to manage their ecosystems. They could introduce competitors of problematic communities or develop mutualistic communities to complement existing ones. Should they do so by building a new community or by adapting an existing one? Building new communities is notoriously difficult, prone to failure, and seems slow compared to the timescales of mutualism and competition (Schweik and English 2012); therefore, adaptation might seem more practicable. However, this study’s findings suggest that communities face structural intertia and may struggle to adapt mutualistically; still, solutions are conceivable. When existing communities spawn new communities that have strong connections with their parents, the new communities can grow faster (Tan 2018). Organizing a spin-off or offshoot community may thus be a viable strategy to fill needed roles in online commuity ecosystems. The “parent” community should create a channel where members can collaborate to design the new community and coordinate their efforts to build it. Of course, doing so is only possible through the efforts of interested members.
@@
A second set of threats emerges from the panel regressions. Although fixed effect panel models are consistent even in the presence of time-constant omitted variables (Wooldridge 2011), time-varying unobserved variables, such as trends in topics’ popularity, can potentially bias estimates. Therefore, readers should not draw causal conclusions from these models. In general, this study reveals the overall trends of competition, mutualism, and overlaps on Reddit at a large scale. However, the wide view and analytical approach obscure the particular mechanisms driving competition and mutualism among these communities and when they are likely to arise. Future fine-grained case studies and experiments should seek to uncover such mechanisms.
@@ Conclusion
Ethical Statement
This work’s mode of inquiry is to analyze large-scale data from social media that is publicly available. Some have concerns about the ethics of such analysis, such as the possibility that the people whose activities created the data we analyze would have been likely to expect their activity would be so used. The author is sympathetic to these serious concerns. That said, the nature of this analysis exposes no individual to scrutiny and the reporting obscures even the identities of the communities analyzed. As a result, the likely and potential harms of this work to individuals are low.
@@ Acknowledgments
Text from a draft of this article was included as part of the author’s PhD dissertation at the Department of Communication at the University of Washington. He thanks his committee: Professors Benjamin Mako Hill, Kirsten Foot, Aaron Shaw, David McDonald and Emma Spiro, for their generous support, wise advice and insightful comments. Versions of this paper received helpful feedback at the International Communication Association’s 2022 annual meeting and at the International Conference for Computational Social Science. Special thanks to Simone Cenci for advice on the regularized S-MAP. Thanks to the Community Data Science Collective for additional feedback and support.